OpenStudy (anonymous):

Test worth 100 points contains 40 questions. There are two point and four point questions on the test. How many of each type?

OpenStudy (anonymous):

x = number of 2 point questions y = number of 4 point questions x + y = 40 2x + 4y = 100 Do you know how to solve for x and y?

OpenStudy (regan):

2x+4y=100 x+y=40

OpenStudy (regan):

sorry

OpenStudy (anonymous):

Thanks.